OpenStudy (anonymous):

I NEED HELP: Find the derivative of the function. F(x) = (x^4 + 3x^2 − 6)^4

OpenStudy (anonymous):

You must use Chain Rule. \[g(x) = x^4 + 3x^2 - 6\]When you have:\[F(x) = f(g(x))\]\[F'(x) = f'(g(x))*g'(x)\]in this case:\[f(x) = x^4\]\[g(x) = x^3 + 3x^2 - 6\]\[f(g(x)) = (x^3 + 3x^2 - 6)^4\] when i do this, i see it as, derivative of whole thing times the derivative of what's inside. \[F'(x) = f'(g(x)) = 4(x^4 + 3x^2 -6)^3 * (4x^3 + 6x)\] usually no need to simplify further

OpenStudy (anonymous):

please ignore second line and typo for what g(x) is. we both know what it really is ^_^

OpenStudy (anonymous):

and the second to last line, after the first equality,lol. but the final answer is right and the initial F'(x) = f'(g(x))*g'(x) is this helpful? any questions?